# Particle identification with machine learning from incomplete data in the ALICE experiment

### v1

@@ 1 Introduction
The particles are identified based on observables derived from signals recorded by the detectors using selection criteria compared with theoretical calculations. The traditional approach, the so-called $\rm{n}_{\sigma}$ method, is to compare the number of standard deviations from the expected value for all detector-specific observables. If a particle has associated PID observables exceeding a certain number of standard deviations, it is rejected. For instance, if one intends to use both TPC and TOF signals, the PID selection would be defined as: $\rm\sqrt{n_{\sigma,TPC}^{2}+n_{\sigma,TOF}^{2}}<\Lambda$, where $\Lambda$ depends on desired balance between purity and efficiency, and typically is in range of 2 to 3. Such an approach is justified when the separation between various particle species is significantly large. However, in reality, the characteristics of different particle species can overlap, and combining information from multiple detectors can become very complex.
@@ 3 Feature Set Embedding and the attention mechanism
The results reported in Ref. [20] clearly show that machine learning algorithms easily outperform the standard method as measured by $F_{1}$ metrics. The proposed attention architecture achieves very good scores for each metric, comparable with other analyzed ML models. At the same time, our model avoids the flaws of other solutions: artificial bias in imputed and case-deleted data and potentially larger complexity of the neural network ensemble.
@@ 4 Domain Adversarial Neural Networks
Naturally, this is a limited solution, which does not allow for full domain alignment of all particle features. Therefore, we will make use of a known machine learning technique called domain adaptation, which aims to learn the discrepancies between two data domains, the labeled source, and the unlabeled target, and translate those to a single hyperspace. The desired classifier is trained and applied to features from the combined latent space. Since the classifier works independently of the initial data domains, it achieves similar performance on both the labeled and unlabeled data (simulated and experimental data in our case).
@@ 5 Conclusions and outlook
Using machine learning for particle identification can result in higher purity and efficiency than standard methods. The current priority is to test PID ML in a real-world analysis of simulation data from the present Run 3 data-taking period. Afterward, we will extend the attention model with domain adaptation and test it on the new real data. Finally, we will be able to achieve regular production of models for Run 3.

### v3

@@ 1 Introduction
The particles are identified based on observables derived from signals recorded by the detectors using selection criteria compared with theoretical calculations such as the Bethe-Bloch formula in case of TPC signals. The traditional approach, the so-called $\rm{n}_{\sigma}$ method, is to compare the number of standard deviations from the expected value for all detector-specific observables. If a particle has associated PID observables exceeding a certain number of standard deviations, it is rejected. For instance, if one intends to use both TPC and TOF signals, the PID selection would be defined as: $\rm\sqrt{n_{\sigma,TPC}^{2}+n_{\sigma,TOF}^{2}}<\Lambda$, where $\Lambda$ depends on desired balance between purity and efficiency, and typically is in range of 2 to 3. Such an approach is justified when the separation between various particle species is significantly large. However, in reality, the characteristics of different particle species can overlap, and combining information from multiple detectors can become very complex.
@@ 2.1 Neural network approach
Nevertheless, it must be considered that it is more difficult to estimate the systematic uncertainties of machine learning models. An example solution is to use dropout, a standard method for reducing overfitting in neural networks, to approximate Bayesian uncertainty as shown in [13].
@@ 3 Feature Set Embedding and the attention mechanism
The results reported in Ref. [21] clearly show that machine learning algorithms easily outperform the standard method as measured by $F_{1}$ metrics. The proposed attention architecture achieves very high scores of $F_{1}$, precision (purity), and recall (efficiency), comparable with other analyzed ML models. At the same time, our model avoids the flaws of other solutions: artificial bias in imputed and case-deleted data and potentially larger complexity of the neural network ensemble.
@@ 4 Domain Adversarial Neural Networks
Naturally, this is a limited solution, which does not allow for full domain alignment of all particle features. Therefore, we will make use of a known machine learning technique called domain adaptation, which aims to learn the discrepancies between two data domains, the labeled source and the unlabeled target, and translate those to a single hyperspace. The desired classifier is trained and applied to features from the combined latent space. Since the classifier works independently of the initial data domains, it achieves similar performance on both the labeled and unlabeled data (simulated and experimental data in our case).
@@ 5 Conclusions and outlook
Using machine learning for particle identification can result in higher purity and efficiency than standard methods. The current priority is to test PID ML in a real-world analysis of simulation data from the present Run 3 data-taking period. Afterward, we will extend the attention model with domain adaptation and test it on the new real data. We will check what levels of disagreement between experimental and Monte Carlo data DANN can handle, and we will estimate model uncertainties. Finally, we will be able to achieve regular production of models for Run 3.